The question is Find the x-intcerpts(if any) for the graph of the quadratic function. F(x)= 2+3x+x^2
Please help me solve the equation.

Answer by stanbon(75887)

Find the x-intercepts(if any) for the graph of the quadratic function.
F(x)= 2+3x+x^2
---
Solve: x^2+3x+2 = 0
Factor:
(x+2)(x+1) = 0
x = -2 or x = -1
====================
Cheers,
Stan H.
============